OpenStudy (anonymous):

PLEASE HELP!!!! A bank account earns 9 percent interest compounded continuously. At what (constant, continuous) rate must a parent deposit money into such an account in order to save 250000 dollars in 9 years for a child's college expenses?

OpenStudy (tkhunny):

You can experiment and gain confidence in your result. i = 0.09 m = Compounding periods per year n = 9 (Total number of years) n*m = Total number of payments j = i/m r = 1+j Formulation \(Pmt\cdot (r + r^{2} + r^{3} + ... + r^{n*m}) = 250000\) \(Pmt\cdot\dfrac{r-r^{n*m+1}}{1-r} = 250000\) Annual m = 1 n*m = 9 g = 0.09 r = 1.09 Pmt = 17614.40414 Semi-Annual m = 2 n*m = 18 g = 0.045 r = 1.045 Pmt = 8908.34869 Quarterly m = 4 Pmt = 4480.49269 Monthly m = 12 Pmt = 1499.48106 Weekly m = 52 Pmt = 346.57201 Daily m = 365 Pmt = 49.39445 Hourly m = 365*24 Pmt = 2.05823 Minutely m = 365*24*60 Pmt = 0.0343039853 Find the continuous formulation and see where it leads.